OpenStudy (anonymous):

Felix exclaims that his quadratic with a discriminant of -1 has no real solutions . Create an equation with a negative discriminate then explain how to find solutions even though they are not real

OpenStudy (anonymous):

Help please quick

OpenStudy (anonymous):

😭😥😥😥😥 please help

OpenStudy (anonymous):

oh im sorry i forgot how to do this maybe @hartnn can help you

OpenStudy (anonymous):

I just need help with the equation

hartnn (hartnn):

discriminant = b^2-4ac choose values of a,b,c such that b^2-4ac = -1

OpenStudy (anonymous):

I got that but how do I make the equation @hartnn

hartnn (hartnn):

assume b= 0, a = 1, so find c from that ? then your equation is \(ax^2+bx+c=0 \\ x^2+c=0\)